# Pair Production in time-dependent Electric field at Finite times

### v4

@@ A.2 Evaluation simple analytical expression for $f(\bm{p},t)$
Using the above relations Eq. (126) and  (127), we approximate the Gauss-hypergeometric functions $f_{1}$ and $f_{2}$ present in the relation of the particle distribution function (see Eq.(59)) as follows:
$f_{1}$ $={}_{2}\mathcal{F}_{1}\left(1+a,1+b;1+c;y\right)$
$+\frac{\Gamma(1+c)\Gamma(1+a+b-c)}{\Gamma(1+a)\Gamma(1+b)}(1-y)^{(c-a-b-1)}{}_{2}\mathcal{F}_{1}\left(c-a,c-b,c-a-b;1-y\right)$
$=\Biggl{(}\frac{c\Gamma(c)\Gamma(c-a-b-1)}{\Gamma(c-a)\Gamma(c-b)}\Biggr{)}\Biggl{(}1+\frac{(a+1)(b+1)}{(2+a+b-c)}(1-y)$
@@
Here, $y$ is a variable that changes with time according to Eq. (47). To compute an approximate expression for the particle distribution function that depends on finite time, we can truncate the power series of the Gauss-hypergeometric functions $f_{1}$and $f_{2}$ given by Eqs. (128) and (129) respectively, up to a specific order. The order of truncation will depend on the desired accuracy and the characteristics of the finite-time behavior under consideration.
Let’s start by approximating the different terms present in the particle distribution relation Eq. (70) and (71), considering only up to the second-order terms and neglecting higher-order terms.
Therefore,
@@
$|\Gamma(\textrm{i}z)|^{2}=\frac{\pi}{z\sinh{(\pi z)}},$ $|\Gamma(1+\textrm{i}z)|^{2}=\frac{\pi z}{\sinh{(\pi z)}},$ $|\Gamma(\frac{1}{2}+\textrm{i}z)|^{2}$ $=\frac{\pi}{\cosh{(\pi z)}}.$ (138)
@@
$f(\bm{p},y)\approx|N^{+}(\bm{p})|^{2}\Bigg{|}\frac{2}{\tau}y(a+b-c)\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln{(1-y)}}+(1-y)\Biggl{[}\Gamma_{1}\Biggl{(}\frac{2}{\tau}yab+\textrm{i}(c-a-b-1)(\omega_{1}-\omega_{0}-\frac{2E_{0}\tau P_{1}}{\omega_{1}})\Biggr{)}$
@@
$1+\frac{P(p_{\parallel},y)}{\omega(\bm{p},y)}$ $\approx\Omega_{0}+(1-y)\Omega_{1}+(1-y)^{2}\Omega_{2}$ (147)
@@
$\Omega_{0}$ $=1+\frac{P_{1}}{\omega_{1}};$ (148)
$\Omega_{1}$ $=\frac{-2E_{0}\tau}{\omega_{1}^{3}}\epsilon^{2}_{\perp}(p_{\perp});$ (149)
@@
$f(\bm{p},y)$ $\approx|N^{+}(\bm{p})|^{2}\Bigl{(}\mathrm{C}_{0}(\bm{p},y)+(1-y)\mathrm{C}_{1}(\bm{p},y)+(1-y)^{2}\mathrm{C}_{2}(\bm{p},y)\Bigr{)}$ (151)
@@
$\mathrm{C}_{0}(\bm{p},y)$ $=4y^{2}\Omega_{0}\omega_{1}^{2}|\Gamma_{2}|^{2}$ (152)
$\mathrm{C}_{1}(\bm{p},y)$ $=4y\Omega_{0}|\Gamma_{1}\overline{\Gamma_{2}}|E_{0}\tau(P_{1}-\omega_{1})\cos{(\Upsilon)}+4y|\Gamma_{2}|^{2}(2\Omega_{0}(\omega_{1}^{2}-\omega_{0}\omega_{1}+(\omega_{1}-P_{1})E_{0}\tau)+\Omega_{1}\omega_{1}^{2}y)$ (153)
@@
$-y\omega_{1}^{2}(10+3\omega_{1}^{2}\tau^{2}-2E_{0}\omega_{1}\tau^{3}+4E_{0}^{2}\tau^{4}))\Bigr{)}\sin{(\Upsilon)}$ (154)
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Bigr{]}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2})$ (155)
Here,
@@
$-4\Omega_{0}\omega_{1}(4+(-\omega_{0}^{2}+\omega_{1}^{2})\tau^{2})y^{2})$ (158)
@@
$-y\omega_{1}^{2}(10+3\omega_{1}^{2}\tau^{2}-2E_{0}\omega_{1}\tau^{3}+4E_{0}^{2}\tau^{4}))$ (159)
A.3 Research Data
This section is included to provide complete clarity on the numerically generated simulation data. As outlined in Section IV.2, we solved the Dirac equation for homogeneous electric fields in Section III to determine the relevant time scales for understanding the dynamics of pair production in a time-dependent electric field.
A.3.1 Longitudinal Momentum spectrum
As discussed in the section on the LMS of created particles, we provide a quantitative estimation to identify the relevant time scales for the dynamics of the longitudinal momentum distribution function, which exhibits distinct behaviors and substructures. The time scales for various values of the electric field parameters are summarized in Tables.
Specifically, one table is provided for each time scale: time scale $t_{cp}$ is analyzed in Table 4, time scale $t_{sep}$ in Table 5, and time scale $t_{dis}$ in Table 6.
         0.2          0.3          0.4
         7          15.272          9.68          6.2
         9          22          12.2          8.3
         11          25.64          15.84          11.19
         13          30.19          18.56          12.03
         15          38.5          21.5          12.05
         17          40          22          12
         19          44          23.4          14
         21          51.3          28.5          15
         23          57.9          33.6          18
Table 4: Numerical data regarding the time scale $t_{cp}$ (Central Peak Formation in distribution function) obtained by evaluating the function $f(p_{\parallel},t)$ for a homogeneous Sauter pulse, $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, at different times. The field strengths $E_{0}$ are given in the first row in terms of the critical field strength $E_{c}$ and the pulse lengths $\tau$ along the first column in units of $[m^{-1}]$.
         0.2          0.3          0.4
         7          18.143          12.63          9.5
         9          25.1          16.92          12.64
         11          31.94          21.09          15.58
         13          38.56          25.36          18.67
         15          45.4          29.65          21.8
         17          51.5          33.5          24.6
         19          57.9          37.53          27.4
         21          64.2          41.8          31
         23          70.6          45.6          33
Table 5: Numerical data regarding the time scale $t_{sep}$ (Peak Separation): At this time, the central peak becomes dominant or the time after the two peaks become distinctly separated. Data obtained by evaluating the function $f(p_{\parallel},t)$ for a homogeneous Sauter pulse, $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, at different times. The field strengths $E_{0}$ are given in the first row in terms of the critical field strength $E_{c}$ and the pulse lengths $\tau$ along the first column in units of $[m^{-1}]$.
         0.2          0.3          0.4
         7          35.143          27.263          24.5
         9          47          34.2          30.11
         11          56.64          40.29          34.88
         13          63.56          47.36          38.56
         15          77.8          58          42.2
         17          80.5          55.5          43.26
         19          88.9          59.52          47.74
         21          97.2          68.9          51.2
         23          112.6          79.8          76
Table 6: Numerical data regarding the time scale $t_{dis}$ (Disappearance of Oscillation): This indicates the time when the oscillations within the central peak fade away or the time after the primary peak (or the left-side peak) ceases to exist. Data obtained by evaluating the function $f(p_{\parallel},t)$ for a homogeneous Sauter pulse, $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, at different times. The field strengths $E_{0}$ are given in the first row in terms of the critical field strength $E_{c}$ and the pulse lengths $\tau$ along the first column in units of $[m^{-1}]$.
A.3.2 Transverse Momentum spectrum
As discussed in the section on TMS of created particles, we provide a quantitative estimation to determine the relevant time scales for the dynamics of the transverse momentum distribution function, which exhibits different behaviors and substructure. We present the time scale for different values of the electric field parameters. It is these data sets and the quality of the fits that have been tabulated here. Specifically, we have generated one table for each time scale; time scale $T_{1}$ is analyzed in Tab.7, time scale $T_{2}$ is analyzed in Tab. 8, and time scale $T_{3}$ is analyzed in Tab. 9, respectively.
0.2 0.3 0.4 0.2 0.3 0.4
7 7.2 2.4 1.98 0.013 0.339 0.075
10 6.9 4.3 2.6 0.048 0.240 0.284
13 6.8 2.6 1.95 0.076 0.273 0.034
16 9 4 1.88 0.179 0.165 0.011
19 6.75 2.3 1.92 0.103 0.464 0.001
Table 7: Numerical data regarding the time scale $T_{1}$ (smooth spectrum begins to reveal an interference pattern, manifesting as a multi-peak structure centered around a non-zero momentum) obtained by evaluating the function $f(p_{\parallel},t)$ for a homogeneous Sauter pulse, $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, at different times. The field strengths $E_{0}$ are given in the first row in terms of the critical field strength $E_{c}$ and the pulse lengths $\tau$ along the first column in units of $[m^{-1}]$. The absolute values obtained are shown on the left and in units of $[m^{-1}]$, as are the relative deviations from our fit $0.28\tau^{0.05}{E_{0}}^{-1.96}$ on the right.
0.2 0.3 0.4 0.2 0.3 0.4
7 15.5 6.8 4.3 0.044 0.141 0.298
10 17.4 6.7 4.5 0.062 0.087 0.162
13 22.7 8.9 4.39 0.041 0.035 0.010
16 21.6 9.3 4.78 0.146 0.050 0.055
19 30 11.1 4.5 0.081 0.021 0.247
Table 8: Numerical data regarding the time scale $T_{2}$ (multi-peak structure in TMS) obtained by evaluating the function $f(p_{\parallel},t)$ for a homogeneous Sauter pulse, $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, at different times. The field strengths $E_{0}$ are given in the first row in terms of the critical field strength $E_{c}$ and the pulse lengths $\tau$ along the first column in units of $[m^{-1}]$. The absolute values obtained are shown on the left and in units of $[m^{-1}]$, as are the relative deviations from our fit $0.11\tau^{0.63}{E_{0}}^{-2.3}$ on the right.
0.2 0.3 0.4 0.2 0.3 0.4
7 26 18 10.2 0.143 0.201 0.156
10 35.4 19.5 10.4 0.002 0.123 0.017
13 43 18 10.2 0.067 0.078 0.137
16 47 15 9.2 0.057 0.430 0.393
19 45 29 14.8 0.069 0.196 0.058
Table 9: Numerical data regarding the time scale $T_{1}$ (small peaks at finite momenta gradually diminish over this time scale in the particle distribution related real particle) obtained by evaluating the function $f(p_{\parallel},t)$ for a homogeneous Sauter pulse, $E\left(t\right)=E_{0}\textrm{sech}^{2}\left(\frac{t}{\tau}\right)$, at different times. The field strengths $E_{0}$ are given in the first row in terms of the critical field strength $E_{c}$ and the pulse lengths $\tau$ along the first column in units of $[m^{-1}]$. The absolute values obtained are shown on the left and in units of $[m^{-1}]$, as are the relative deviations from our fit $0.66\tau^{0.49}{E_{0}}^{-1.79}$ on the right.
@@ A.3.2 Transverse Momentum spectrum
A.4 Fourier analysis
To analyze the transverse momentum spectrum at finite time using the Fourier analysis, we decompose the time-dependent electric field $E(t)$ into cosine and sine components.
@@ A.4 Fourier analysis
$E(t)$ $=\sum_{i}\tilde{E}(\nu_{i})e^{i\nu_{i}t}$ (162)
We calculate the frequency $\nu$ using the Eq. (161) at some time, and the total probability can be compute by absolute sum of different amplitude of pair production due to different $\nu_{i}$ at the finite time $t=t_{\text{out}}$ and plot the total probability as the function of $p_{\perp}$ as shown in Fig.20.

### v5

@@ A.1 Evaluation approximate analytical expression for the distribution function
Here, $y$ is transformed time variable.
To derive an approximate expression for the distribution function, we truncate the power series of the Gauss-hypergeometric functions $f_{1}$ and $f_{2}$, as presented in Eqs. (142) and (143), respectively, up to a specific order. In doing so, we retain only the second-order terms in the expressions for the distribution function, as given by Eq. (142) and (143), and neglect higher-order terms of order $\mathcal{O}((1-y)^{2})$.
@@
$\Gamma(\textrm{i}z)\overline{\Gamma(\textrm{i}z)}=\frac{\pi}{z\sinh{(\pi z)}},$ $|\Gamma(1+\textrm{i}z)|^{2}=\frac{\pi z}{\sinh{(\pi z)}},$ $|\Gamma(\frac{1}{2}+\textrm{i}z)|^{2}$ $=\frac{\pi}{\cosh{(\pi z)}}.$ (152)
where $\overline{\Gamma(z)}$ is complex conjugate of the $\Gamma(z).$
@@
$f(\bm{p},y)\approx|N^{(+)}(\bm{p})|^{2}\Bigg{|}\frac{2}{\tau}y(a+b-c)\Gamma_{2}e^{-\textrm{i}\tau\omega_{1}\ln{(1-y)}}+(1-y)\Biggl{[}\Gamma_{1}\Biggl{(}\frac{2}{\tau}yab+\textrm{i}(c-a-b-1)(\omega_{1}-\omega_{0}-\frac{2E_{0}\tau P_{1}}{\omega_{1}})\Biggr{)}$
@@
$1+\frac{P(p_{\parallel},y)}{\omega(\bm{p},y)}$ $\approx\Omega_{0}+(1-y)\Omega_{1}+(1-y)^{2}\Omega_{2},$ (161)
@@
$\Omega_{0}$ $=1+\frac{P_{1}}{\omega_{1}},$ (162)
$\Omega_{1}$ $=\frac{-2E_{0}\tau}{\omega_{1}^{3}}\epsilon^{2}_{\perp}(p_{\perp}),$ (163)
@@
$f(\bm{p},y)$ $\approx|N^{(+)}(\bm{p})|^{2}\Bigl{(}\mathrm{C}_{0}(\bm{p},y)+(1-y)\mathrm{C}_{1}(\bm{p},y)+(1-y)^{2}\mathrm{C}_{2}(\bm{p},y)\Bigr{)},$ (165)
@@
$\mathrm{C}_{0}(\bm{p},y)$ $=4y^{2}\Omega_{0}\omega_{1}^{2}|\Gamma_{2}|^{2},$ (166)
@@
$-y\omega_{1}^{2}(10+3\omega_{1}^{2}\tau^{2}-2E_{0}\omega_{1}\tau^{3}+4E_{0}^{2}\tau^{4}))\Bigr{)}\sin{(\Upsilon)}.$ (168)
Let us denote
@@
$-4\Omega_{0}\omega_{1}(4+(-\omega_{0}^{2}+\omega_{1}^{2})\tau^{2})y^{2}),$ (171)
@@
$-y\omega_{1}^{2}(10+3\omega_{1}^{2}\tau^{2}-2E_{0}\omega_{1}\tau^{3}+4E_{0}^{2}\tau^{4})).$ (172)
Therefore,
$+4\Omega_{2}y^{2}\omega_{1}^{2}+\frac{\Omega_{0}}{4\omega_{1}^{2}(1+\tau^{2}\omega_{1}^{2})}\sigma_{0}\Bigr{]}+\frac{|\Gamma_{1}\overline{\Gamma_{2}}|}{2\omega_{1}^{2}(4+\omega_{1}^{2}\tau^{2})}(\cos{(\Upsilon)}\sigma_{1}+\sin{(\Upsilon)}\sigma_{2}).$ (173)
@@
A.2 Fourier analysis
We decompose the time-dependent electric field $E(t)$ into cosine and sine components to analyze the TMS at finite time using the Fourier analysis.
@@ A.2 Fourier analysis
$E(t)$ $=\sum_{-n}^{n}\tilde{E}(\nu_{n})e^{i\nu_{n}t}$ (176)
We calculate the frequency $\nu$ using the Eq. (175) at time $t$, and the total probability can be compute by absolute sum of different amplitude of pair production due to different $\nu_{n}$ at the finite time $t=t_{\text{out}}$ and plot the total probability as the function of $p_{\perp}$ as shown in Fig.20.